# Divide et Impera: Learning impedance families for peg-in-hole assembly

### v1

@@ Abstract
This paper addresses robotic peg-in-hole assembly using the framework of Elementary Dynamic Actions (EDA). Inspired by motor primitives in neuromotor control research, the method leverages three primitives: submovements, oscillations, and mechanical impedances (e.g., stiffness and damping), combined via a Norton equivalent network model. By focusing on impedance parameterization, we explore the adaptability of EDA in contact-rich tasks. Experimental results, conducted on a real robot setup with four different peg types, demonstrated a range of successful impedance parameters, challenging conventional methods that seek optimal parameters. We analyze our data in a lower-dimensional solution space. Clustering analysis shows the possibility to identify different individual strategies for each single peg, as well as common strategies across all pegs. A neural network model, trained on the experimental data, accurately predicted successful impedance parameters across all pegs. The practical utility of this work is enhanced by a success-predictor model and the public availability of all code and CAD files. These findings highlight the flexibility and robustness of EDA; show multiple equally-successful strategies for contact-rich manipulation; and offer valuable insights and tools for robotic assembly programming.
@@ I Introduction
Lately, robotics has seen promising developments in contact-rich manipulation, such as model-based methods [7, 8, 5] and learning-based approaches [9, 3, 10, 6]. Model-based methods leverage mathematical representations of contact dynamics, ensuring that the robots can predict and react to contact forces accurately. Learning-based approaches utilize probabilistic models to handle uncertainties in contact interactions. Learning approaches, particularly those based on reinforcement learning and neural networks, enable robots to learn optimal contact strategies through trial and error, leading to adaptable manipulation skills.
@@
Sampling-based optimization methods have also gained traction, particularly in scenarios where gradient information is not readily available. Most of these sampling-based methods are Zeroth-order methods, which do not rely on gradients and sample multiple actions and update control parameters based on outcomes without needing derivative information [15]. Variations such as STOMP [16], PI² [17, 18], and MPPI [19, 20, 21] have become popular. First-order methods, involving randomized smoothing for gradient-based optimization, are also employed [22, 5]. However, these methods are often constrained by the choice of the cost function, which will only output one set of control parameters, even though multiple solutions might be feasible. Moreover, those methods often lead to trivial solutions that merely resemble position controllers [17, 23].
@@
Other common approaches in robotic control are Dynamic Movement Primitives (DMP) and Elementary Dynamic Actions (EDA). Both approaches are inspired by human motor control research [24, 25, 26, 27] and try to leverage insights into how humans achieve complex, coordinated movements [28, 29]. It is hypothesized that keys to human performance are building blocks of coordination, as originally proposed by Sherrington [30]. These building blocks are conventionally called “motor primitives” and are basic motor actions that can generate a rich repertoire of behaviors. Bernstein further expanded on this concept by introducing the idea of “repetition without repetition” [31]. He observed that humans can accomplish tasks in numerous ways, adapting to subtle environmental changes and learning strategies for movement. This adaptability is a key ability of the human motor system. Latash argued that the motor system does not operate with explicit costs but rather utilizes a subspace of multiple solutions, enabling flexible and effective movement strategies [32]. Wolpert et al. emphasized the importance of motor learning, highlighting that feed-forward control is essential for natural movements, as feedback mechanisms can be too slow[33, 34].
The main motivation of this work is to investigate whether we can achieve adaptability comparable to that observed in human motor control studies by using a “human motor control inspired” robot controller.
In this paper, our main method will be EDA, originally introduced by [26, 27]. The key to this approach is breaking down complex physical interaction tasks into simpler, fundamental actions that can be more easily controlled and adapted. EDA consists of three distinct classes of motor primitives: submovements and oscillations as kinematic primitives [35], and mechanical impedances as distinct primitives to manage physical interaction [36, 26, 37]. The three primitives can be combined using a Norton equivalent network model [38]. More details about the primitives of EDA and the Norton equivalent network model can be found in appendix A.
While EDA has demonstrated remarkable success in contact-rich manipulation [39, 27, 40], parameterizing the distinct primitives remains a key challenge. This paper aims to elucidate the choices of parameters for interaction primitives, specifically impedance (e.g., stiffness and damping). Our use-case will focus on peg-in-hole assembly which is a common task in robotic research and industrial assembly (see fig. 1). To our knowledge, this paper is the first to go beyond generating a controller for contact-rich manipulation, but aims to illuminate the structure, effects, and distribution of successful parameter choices for peg-in-hole assembly. Our goal is to provide novel insights that can enhance our understanding and application of EDA in practical scenarios, as well as to help in finding more intuitive solutions to complex manipulation tasks.
@@
Figure 1: Exemplary peg-in-hole assembly using a KUKA LBR robot. All control parameters were defined with respect to the spatially fixed coordinate frame, highlighted in orange.
The main contributions of this paper are to investigate four distinct claims. The knowledge gained from this investigation will culminate in a success predictor for robotic peg-in-hole assembly, which is intended to help less-experienced programmers to use our method.
  • Claim 1: We demonstrate that Elementary Dynamic Actions (EDA) do not restrict successful assembly to a single optimal set of parameters but instead enable a range of successful impedance parameters. This finding challenges the conventional method of finding a single optimal solution in robotic assembly tasks.
  • Claim 2: Our analysis reveals the existence of a non-arbitrary, lower-dimensional representation of successful assembly strategies. This simplification is achieved through Principal Component Analysis (PCA), which shrinks the parameter space dimension from eight to three without substantially removing information.
  • Claim 3: We identify that different assembly strategies are effective for individual peg types. This highlights the adaptability of EDA to accommodate variations and provide a tailored approach to different assembly challenges.
  • Claim 4: Across various pegs, common assembly strategies emerge, suggesting that impedance settings may be generalized across different tasks. This would support the development of more streamlined and efficient programming practices.
  • Application: We introduce a neural network model as a success-predictor for peg-in-hole assembly. This model offers a practical tool to predict assembly outcomes based on impedance parameters and can serve as a valuable resource for less-experienced robot programmers.
@@ II Methods
This section outlines the theoretical framework and experimental protocol used in this work to investigate the role of impedance parameter choices in robotic peg-in-hole assembly. We employed a systematic approach to evaluate the role of EDA and their components to achieve robotic assembly operations. All methods introduced here were used to investigate the main claims of the paper.
II-A Assembly strategy for peg-in-hole tasks
@@
For our assembly application, we used a KUKA LBR robot and the Cartesian Impedance Controller of the KUKA Sunrise framework 1.17. Details of the algorithms can be found in [41]. The robot application code is provided on Github [^1] https://github.com/mosesnah-shared/ImpedanceClassification/tree/main/Sunrise.
@@
We programmed an application to assemble four different peg-in-hole assemblies: square, triangular, hexagonal, and cylindrical pegs. Between the workpieces and the respective insertion holes was a nominal clearance of 0.14 mm (square peg), 0.18 mm (hexagonal peg), and 0.20 mm (circular and triangular pegs). All pieces were printed with a PRUSA i3 MK3 3D-printer using PLA filament. All components except the pegs were printed with $20\%$ infill with gyroid pattern. Pegs were printed with higher infill ($>70\%$) to guarantee structural strength. The CAD data of all pieces can be found on Dropbox [^2] https://www.dropbox.com/scl/fo/ljrkha4bh17v4scy8uy9l/AL4zqjOkfwap6kvq8QBylWQ?rlkey=qqyxf8euipigzhlokfo5x2nr6&dl=0.
@@
Algorithm 1 shows the structure of the assembly application. For our application, we used the torque sensors of the robot to measure external contact (Algorithm 1, line 3-5 and 14-20).
@@
Algorithm 1 Primitives for peg-in-hole assembly: submovements (orange), oscillations (blue), and mechanical impedance (green). The coordinates related to the algorithm can be seen in fig. 1. The impedance parameters in line 11-13 are exemplary and were varied throughout the experiments. The frame-based discrete movements and oscillations remained fixed throughout the experiments.
@@
The overall assembly strategy was inspired by observations of humans performing assembly tasks, often without visual feedback. Humans frequently rely on tactile feedback and subtle movements, such as wiggling the workpiece, to align and insert components accurately. All primitives of EDA were used. For the kinematic primitives (submovements and oscillations), the frame-based programming procedure of KUKA Sunrise was used. For the impedance parameters, the diagonal terms of the translational and rotational stiffness matrices and the translational and rotational damping ratios were varied throughout the experiments.
@@ II-B Experimental protocol
Finding feasible impedance values is not trivial. Our intention was to test combinations of impedance values for peg-in-hole assembly. Due to the curse of dimensionality [43], it was not practical to test all possible combinations of impedance values. Hence, we assigned three different categories for each of the parameter ranges: small, medium, and large. The exact values of the three categories can be found in appendix B.
@@
After the experiments, we investigated four different claims:
  • Claim 1: Success with a range of parameters
  • Claim 2: Lower dimensional representation of solutions
  • Claim 3: Different assembly strategies for individual pegs
  • Claim 4: Common assembly strategies across all pegs
The methods used to evaluate these claims are explained in detail in the following.
II-C Claim 1: Success with a range of parameters
II-C1 Success Rate
@@ II-C2 Linear regression model
To investigate Claim 1—that EDA enables a range of successful parameters, rather than one optimal set—we used linear regression models to fit the success rate (dependent variable) as a function of the impedance parameters (independent variables) for each peg. The goodness-of-fit for each model was assessed using the coefficient of determination ($R^{2}$), which measures the variance in the dependent variable that is predictable from the independent variable. The statistical significance of the regression slope was tested against the null hypothesis of no relation between the success rate and the impedance parameters examined. The results were interpreted to understand how different impedance parameter choices affected the outcome of the peg-in-hole task. For example, a significant positive slope in the relationship between success rate and the stiffness $k_{z}$ implied that higher vertical stiffness tended to improve task success.
II-D Claim 2: Lower dimensional representation of solutions
II-D1 Principal Component Analysis
Principal Component Analysis (PCA) is a statistical technique used to reduce the complexity of high-dimensional data by projecting it onto a smaller number of dimensions, known as principal components (PCs). This technique is particularly useful for revealing hidden patterns and explaining variance in data.
@@ II-D1 Principal Component Analysis
For our study, we applied PCA to the impedance solutions obtained from our peg-insertion task to test if a non-arbitrary lower-dimensional subspace could account for our data (Claim 2). We projected the data into a subspace defined by the PCs that captured the most variance. The main PCs provide insight into the principal directions in the data. Moreover, a significant reduction in dimension, from eight to three or fewer, would enable us to visualize the 8-dimensional solution space in the lower-dimensional PC space. Specifically, to extract PCs from the impedance solutions, we utilized the pca function from MATLAB 2022b.
As PCA is data-driven, it may identify a lower degree-of-freedom space in random data. To ensure that any reduction of the solution space observed here was not a random artifact, we compared the slope of the variance-accounted-for by the first three dominant PCs to the slope of the variance-accounted-for by an equally-sized random subset of trials. We conducted a pair-wise t-test between slopes of the variance-accounted-for in the successful trials and the random trials. If the slopes in the successful trials were significantly steeper (i.e., the probability that the slopes were the same was $\alpha\leq 0.05$), we concluded that the PCs extracted from the successful trials were not a random artifact. Specifically, we could claim that there existed a non-arbitrary impedance solution subspace.
@@
Additionally, we projected both successful and unsuccessful trials into the identified PC subspace to visually assess their distribution. If unsuccessful trials consistently fell outside the identified subspace, this would further support the non-arbitrary nature of the impedance solution subspace.
II-E Claim 3: Different assembly strategies for individual pegs
II-E1 K-means Clustering
@@
To investigate Claim 3—that families of solutions encode different assembly strategies—-we utilized k-means clustering on successful trials within our 8-dimensional impedance solution space. The identified clusters would represent distinct common impedance solution strategies. We employed MATLAB 2022b’s k-means clustering algorithm with an online phase update and 5 replicates [44, 45]. The online updates, wherein each sample was individually reassigned to a new cluster if it reduced the sum of point-to-centroid distances, ensured convergence to a local minimum. The use of multiple replicates added robustness to the clustering process.
To identify the effective number of clusters we computed the silhouette value $s_{i}$ of each data point using the equation
$s_{i}=\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{max(a_{i},b_{i})},$ (2)
where $i$ is the $i-$th data point, $a_{i}$ is the average distance between the $i-$th point and all other data points within the cluster, and $b_{i}$ is the avg distance between the $i-$th point and all other data points within the next nearest cluster. From the above definition it is clear that $-1\leq s_{i}\leq 1$. Values close to 1 indicate well-clustered data, negative values suggest misclustered data, and values near zero indicate the datum is near the border of two clusters [46, 47].
With this, the overall clustering effectiveness was evaluated by the silhouette score, $SC$, defined as:
$SC=\max_{k}s(k),$ (3)
It is important to note that k-means clustering must have at least two clusters and can have at most as many clusters as the number of data points provided. In the latter case, the silhouette coefficient will approach 1 as each point provides its own cluster. Visual representation was facilitated by projecting the impedance solution data for each peg onto the PCA space established previously.
We conducted a one-way MANOVA to determine if the identified clusters were distinct. Detailed methodology can be found in appendix C.
II-F Claim 4: Common assembly strategies across all pegs
II-F1 Neural Networks for Binary Classification

### v2

# Divide et Impera: Decoding Impedance Strategies for Robotic Peg-in-Hole Assembly

@@ Abstract
This paper investigates robotic peg-in-hole assembly using the Elementary Dynamic Actions (EDA) framework, which models contact-rich tasks through a combination of submovements, oscillations, and mechanical impedance. Rather than focusing on a single optimal parameter set, we analyze the distribution and structure of multiple successful impedance solutions, revealing patterns that guide impedance selection in contact-rich robotic manipulation. Experiments with a real robot and four different peg types demonstrate the presence of task-specific and generalized assembly strategies, identified through K-means Clustering. Principal Component Analysis (PCA) is used to represent these findings, highlighting patterns in successful impedance selections. Additionally, a neural-network-based success predictor accurately estimates feasible impedance parameters, reducing the need for extensive trial-and-error tuning. By providing publicly available code, CAD files, and a trained model, this work enhances the accessibility of impedance control and offers a structured approach to programming robotic assembly tasks, particularly for less-experienced users.
@@ I Introduction
Lately, robotics has seen promising developments in contact-rich manipulation, such as model-based methods [7, 8, 5] and learning-based approaches [9, 3, 10, 6]. Model-based methods leverage mathematical representations of contact dynamics, ensuring that the robots can predict and react to contact forces accurately. Learning-based methods leverage probabilistic models to manage uncertainties in contact interactions. In particular, approaches such as reinforcement learning and neural networks allow robots to develop optimal contact strategies through trial and error, fostering adaptable manipulation skills.
@@
Sampling-based optimization methods have also gained traction, particularly in scenarios where gradient information is not readily available. Most of these sampling-based methods are Zeroth-order methods, which do not rely on gradients and sample multiple actions and update control parameters based on outcomes without needing derivative information [15]. Variations such as STOMP [16], PI² [17, 18], and MPPI [19, 20, 21] have become popular. First-order methods, involving randomized smoothing for gradient-based optimization, are also employed [22, 23, 5]. The approaches stochastically smooth the underlying landscape of the objective function to enhance robust planning and control [24, 25, 26].
The methods discussed above often rely on predefined cost functions. These functions aim to identify a single “optimal” set of control parameters. They frequently converge to trivial solutions, such as uniformly increasing all parameters, resulting in behavior resembling position controllers [17, 27]. While multiple feasible solutions are common, their distribution and interdependence have note been not investigated in these approaches.
@@
Other approaches in contact-rich robotic control are Dynamic Movement Primitives (DMP) and Elementary Dynamic Actions (EDA). Both approaches are inspired by human motor control research [28, 29, 30, 31] and try to leverage insights into how humans achieve complex, coordinated movements [32, 33]. It is hypothesized that keys to human performance are building blocks of coordination, as originally proposed by Sherrington [34]. These building blocks are conventionally called “motor primitives” and are basic motor actions that can generate a rich repertoire of behaviors. Bernstein further expanded on this concept by introducing the idea of “repetition without repetition” [35]. He observed that humans can accomplish tasks in numerous ways, adapting to subtle environmental changes and learning strategies for movement. This adaptability is a key ability of the human motor system. Latash argued that the motor system does not operate with explicit costs but rather utilizes a subspace of multiple solutions, enabling flexible and effective movement strategies [36]. Wolpert et al. emphasized the importance of motor learning, highlighting that feed-forward control is essential for natural movements, as feedback mechanisms can be too slow[37, 38].
The main motivation of the work presented here is to demonstrate that contact-rich robotic tasks generally do not have a single optimal solution; rather, multiple parameterizations can successfully achieve the desired performance. Drawing inspiration from principles well-established in human motor control, we propose that there exists a set of viable solutions capable of adequately performing a given task. To explore this, we adopt a robot control framework based on Elementary Dynamic Actions (EDA) [30, 39, 31, 40], aiming to achieve adaptability comparable to human motor performance. Although EDA has demonstrated notable success in contact-rich manipulation tasks [41, 39], effectively parameterizing its dynamic primitives—particularly stiffness and damping—remains a key challenge.
This paper examines peg-in-hole assembly, a foundational task in robotics (see Fig. 1). Peg-in-hole assembly serves as a representative application for numerous industrial robotic tasks, showcasing essential aspects of contact-rich manipulation. Examples include assembling engine components in the automotive industry, connecting circuit board plugs in electronics manufacturing, and future human-centric applications like furniture assembly.
We present a comprehensive investigation into the influence of impedance parameters, such as stiffness and damping, on task success in robotic peg-in-hole assembly. Our contributions are:
  1. 1.
    Through Principal Component Analysis (PCA) and K-means Clustering, we reveal two key insights: (a) task-specific strategies tailored to individual peg types and (b) generalized strategies applicable across various peg shapes.
  2. 2.
    We develop a neural network model to predict successful impedance parameters, providing a practical tool to support robotic programmers, particularly to aid those with limited experience.
@@
Figure 1: Exemplary peg-in-hole assembly using a KUKA LBR robot. All control parameters were defined with respect to the spatially fixed coordinate frame highlighted in orange.
@@ II Methods
This section outlines the theoretical framework, experimental protocol, and computational techniques employed to investigate the role of impedance parameters in robotic peg-in-hole assembly. Throughout the remainder of the paper, we consider a system with $n$ degrees of freedom (DOF).
II-A Impedance Control
Impedance control [42, 43] models the robot controller as a Norton equivalent network [44], defining the relationship between generalized force output $\mathbf{F}\in se^{\star}(3)$ and motion input $\dot{\mathbf{x}}\in se(3)$ at the interaction port $(\mathbf{F},\dot{\mathbf{x}})$ (Fig. 2B). The controller generates interaction forces $\mathbf{F}$ by coupling the difference between a virtual trajectory $\mathbf{x}_{0}\in SE(3)$ and the current robot pose $\mathbf{x}\in SE(3)$ with mechanical impedance $\mathbf{Z}:se(3)\rightarrow se^{\star}(3)$. Modulating $\mathbf{Z}$ allows control over the dynamics of physical interaction. For example, tactile exploration and manipulation of delicate objects benefit from low stiffness, while tasks like drilling demand high stiffness [45]. Under the assumption that the environment is an admittance, mechanical impedance can be linearly superimposed even though each mechanical impedance is a nonlinear operator [43, 46]. For $i$-number of impedances: $\mathbf{Z}=\sum\mathbf{Z}_{i}$. More information about impedance control can be found in [42, 43, 47].
II-B Elementary Dynamic Actions (EDA)
EDA relates impedance control to principles of human motor control by defining the virtual trajectory $\mathbf{x}_{0}$ through two key motor primitives: discrete movements and oscillatory motions [40, 31]. Additionally, it incorporates mechanical impedance as a distinct dynamic primitive to manage physical interaction [46, 30, 48]. These three primitives are integrated within a Norton equivalent network model (Fig. 2B).
Figure 2: (A) The three primitives of Elementary Dynamic Actions (EDA). Submovements and oscillations are kinematic primitives used to generate discrete and periodic movements respectively [40], and mechanical impedance is a dynamic primitive used to handle physical interaction [46, 45, 49, 50, 51]. (B) The three primitives are combined using a Norton equivalent network model [44].
II-B1 Submovement
A submovement $\mathbf{x}_{0}$ is a smooth trajectory in which its time derivative is a unimodal function, i.e., has a single peak value:
$\dot{\mathbf{x}}_{0}(t)=\mathbf{v}\ \hat{\sigma}(t)$ (1)
II-B2 Oscillation
An oscillation $\mathbf{x}_{0}$ is a smooth non-zero trajectory which is a periodic function:
$\forall t>0:~{}~{}\exists T>0:~{}~{}\mathbf{x}_{0}(t)=\mathbf{x}_{0}(t+T)$ (2)
II-B3 Mechanical Impedance
Mechanical impedance serves as a dynamic primitive in the EDA framework to regulate physical interactions between the robot and its environment. As discussed in Sec. II-A, mechanical impedance $\mathbf{Z}$ defines the relationship between the displacement $\Delta\mathbf{x}$ (difference between virtual and actual trajectories) and the resulting force $\mathbf{F}$. An overview of the literature on mechanical impedance in human motor control can be found in [53].
II-C Assembly strategy for peg-in-hole tasks
@@
For our assembly application, we used a KUKA LBR robot and a Cartesian Impedance Controller [47]. The robot application code is provided on Github [^1] https://github.com/mosesnah-shared/ImpedanceClassification/tree/main/Sunrise.
@@
We programmed an application to assemble four different peg-in-hole types: square, triangular, hexagonal, and cylindrical pegs. Between the workpieces and the respective insertion holes was a nominal clearance of 0.14 mm (square peg), 0.18 mm (hexagonal peg), and 0.20 mm (circular and triangular pegs). All pieces were printed with a PRUSA i3 MK3 3D-printer using PLA filament. All components except the pegs were printed with $20\%$ infill with gyroid pattern. Pegs were printed with higher infill ($>70\%$) to guarantee structural strength. The CAD data of all pieces can be found on Dropbox [^2] https://www.dropbox.com/scl/fo/ljrkha4bh17v4scy8uy9l/AL4zqjOkfwap6kvq8QBylWQ?rlkey=qqyxf8euipigzhlokfo5x2nr6&dl=0.
@@
Algorithm 1 shows the structure of the assembly application. For our application, we used the torque sensors of the robot to detect external contact (Algorithm 1, line 3-5 and 14-20).
@@
Algorithm 1 Primitives for peg-in-hole assembly: submovements (orange), oscillations (blue), and mechanical impedances (green). The coordinates related to the algorithm can be seen in Fig. 1. The impedance parameters in line 11-13 are exemplary and were varied throughout the experiments. The frame-based discrete movements and oscillations remained fixed throughout the experiments.
@@
The overall assembly strategy was inspired by observations of humans performing assembly tasks, often without visual feedback. Humans frequently rely on tactile feedback and subtle movements, such as wiggling the workpiece, to align and insert components accurately.
For the peg-in-hole assembly, we chose a fixed strategy for the zero-force trajectory, $\mathbf{x}_{0}$, and investigated the influence of different choices of impedance parameters, $\mathbf{Z}$. For the kinematic primitives (submovements and oscillations), the frame-based programming procedure of KUKA Sunrise was used. This design choice allowed us to focus our analysis solely on the solution space of impedance parameters, excluding variations in the kinematic properties of the zero-force trajectory. Notably, prior work has explored the inverse approach, i.e., varying the kinematic primitives with constant dynamic primitives [54, 55]. For the impedance parameters, the diagonal terms of the translational and rotational stiffness matrices and the translational and rotational damping ratios were varied throughout the experiments.
@@ II-D Experimental protocol
Finding feasible impedance values is not trivial. Our intention was to test combinations of impedance values for peg-in-hole assembly. Due to the curse of dimensionality [57], it was not practical to test all possible combinations of impedance values. Hence, we assigned three different categories for each of the parameter ranges: small, medium, and large. The exact values of the three categories can be found in Appendix A.
@@
II-E Computational methods
After conducting experiments, we investigated key aspects of successful peg-in-hole assembly:
  • The success rates associated with different impedance parameter combinations.
  • The variability of assembly strategies tailored to individual peg types and their low-dimensional representation.
  • The identification of common strategies, applicable across various peg shapes, represented in a low-dimensional space.
The methods used to derive these findings are detailed in the following subsections.
II-E1 Linear regression to fit the success rate
@@
To understand how different impedance parameter choices affected the outcome of the peg-in-hole task, we used linear regression to fit the success rate (dependent variable) as a function of the impedance parameters (independent variables) for each peg. A linear model was fit to each success rate distribution for each impedance parameter and peg, and the goodness-of-fit was assessed using the coefficient of determination ($R^{2}$), which measures the variance in the dependent variable that is predictable from the independent variable. The resulting linear fit was statistically tested against the degenerate null slope, testing the null hypothesis of no relationship between the success rate and the impedance parameters. For example, a significant positive slope in the relationship between success rate and the stiffness $k_{z}$ implied that higher vertical stiffness tended to improve task success.
II-E2 PCA to identify a low dimensional representation
To reduce the complexity of high-dimensional data and reveal (potential) hidden patterns in data, we used PCA to project the data onto a lower dimensional subspace, often referred to as Principal Components (PCs). The data were pre-processed using the algorithm proposed by West et al. [58], i.e., the input dataset was centered with respect to its mean thus removing an otherwise implicit bias in the first principal component. Moreover, a significant reduction in dimension, from eight to three or fewer, would enable us to visualize the 8-dimensional solution space in the lower-dimensional PC space.
@@ II-E2 PCA to identify a low dimensional representation
We applied PCA to the impedance solutions obtained from our peg-insertion task to test if a non-arbitrary lower-dimensional subspace could account for our data. We projected the data into a subspace defined by the PCs that captured the most variance. The main PCs provide insight into the principal directions in the data. To extract PCs from the impedance solutions, we utilized the pca function from MATLAB 2022b.
To verify that the observed solution space reduction was not random, we compared the variance explained by the first three dominant PCs to that of an equally-sized random subset of trials. More detailed information about this statistical test can be found in Appendix B.
@@
Finally, we projected both successful and unsuccessful trials into the identified PC subspace to visually assess their distribution. If unsuccessful trials consistently fell outside the identified subspace, this would further support the non-arbitrary nature of the impedance solution subspace.
II-E3 K-means Clustering to reveal different assembly strategies
@@
We utilized K-means Clustering on successful trials within our eight-dimensional impedance solution space to investigate clusters that represent distinct common impedance solution strategies. We employed MATLAB 2022b’s K-means Clustering algorithm with an online phase update and five replicates [59, 60]. The online updates, wherein each sample was individually reassigned to a new cluster if it reduced the sum of point-to-centroid distances, ensured convergence to a local minimum. The use of multiple replicates added robustness to the clustering process. More details about the clustering process can be found in Appendix D.
Since PCA-based dimensionality reduction removes variance that may contain important structure within the impedance space, clustering was performed on the original eight-dimensional space to fully capture all impedance parameter relationships. However, for better visualization, we projected the resulting clusters onto a 2D or 3D PCA-reduced space, allowing for an intuitive representation while maintaining the integrity of the clustering process.
To assess the distinctiveness of the identified clusters, we conducted a one-way MANOVA. Details of this statistical validation can be found in Appendix E.
II-E4 Neural network to reveal common assembly strategies